how many permutations can be made from selecting 4 college courses out of 10

and i had some trouble understanding,
four students can choose to write on one of 5 essay topics for their english class .
how many different essays can the student write without repetition of the topics?
how many different essays can the student write with repetition of the topics?

Answer by stanbon(75887)

how many permutations can be made from selecting 4 college courses out of 10
10P4 = 10*9*8*7 = 5040
-------------------------------
four students can choose to write on one of 5 essay topics for their english class. how many different essays can the students write without repetition of the topics?
--
Ans: 5*4*3*2 = 120
The assumption here is that order makes a difference.
-------
how many different essays can the students write with repetition of the topics?
5^4 = 625
===============
Cheers,
Stan H.
===========